# Two-grid convergence theory for symmetric positive semidefinite linear systems

### v1

@@ 3. Convergence analysis of Algorithm 1
Define
$r:=\operatorname*{rank}(A)\quad\text{and}\quad s:=\operatorname*{rank}(A_{\rm c}).$
@@
which, together with (2.8), yields $\|E_{\rm TG}\|_{A}=0$. Such an extreme case seldom occurs in practice. In what follows, we only consider the nontrivial case $s<r$.
@@
Under the assumptions of Algorithm 1 and the condition (2.10), one has
@@
where $F_{\rm TG}$ is given by (2.9).
@@
$\|E_{\rm TG}\|_{A}^{2}$ $=1-\min_{\mathbf{v}\in\mathcal{R}(A)\backslash\{0\}}\frac{\mathbf{v}^{T}F_{\rm TG}\mathbf{v}}{\mathbf{v}^{T}\mathbf{v}}$
@@
$=1-\min_{\mathbf{v}_{r}\in\mathbb{R}^{r}\backslash\{0\}}\frac{\mathbf{v}_{r}^{T}\Lambda_{r}\mathbf{v}_{r}}{\mathbf{v}_{r}^{T}\mathbf{v}_{r}}$
@@
$\|E_{\rm STG}\|_{A}^{2}$ $=\max_{\mathbf{v}\in\mathcal{R}(A)\backslash\{0\}}\frac{\mathbf{v}^{T}(I-F_{\rm TG})^{2}\mathbf{v}}{\mathbf{v}^{T}\mathbf{v}}$
@@
$=\max_{\mathbf{v}_{r}\in\mathbb{R}^{r}\backslash\{0\}}\frac{\mathbf{v}_{r}^{T}(I_{r}-\Lambda_{r})^{2}\mathbf{v}_{r}}{\mathbf{v}_{r}^{T}\mathbf{v}_{r}}.$
@@
Based on Lemma 3.1, we can derive the following convergence identity, which is the main result of this section.
@@
Under the assumptions of Lemma 3.1, it holds that
(3.7) $\|E_{\rm TG}\|_{A}=\sqrt{1-\sigma_{\rm TG}},$
where
(3.8) $\sigma_{\rm TG}=\lambda_{n-r+s+1}\big{(}\widetilde{M}A(I-\Pi_{A})\big{)}.$
@@
According to the proof of Lemma 3.1, we deduce that $A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}(I-\Pi)+\Pi$ has $r$ positive eigenvalues and eigenvalue 00 with multiplicity $n-r$.
@@
$A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}(I-\Pi)+\Pi=Q_{2}\begin{pmatrix}I_{s}&X_{2}\\ 0&X_{3}\end{pmatrix}Q_{2}^{T},$
from which we deduce that $X_{3}$ has $r-s$ positive eigenvalues and eigenvalue 00 with multiplicity $n-r$.
@@
The identity (3.7) provides a direct way to analyze the convergence properties of Algorithm 1. A two-sided estimate for $\|E_{\rm TG}\|_{A}$ can be obtained by applying (3.7) and the following eigenvalue inequality; see, e.g., [12, Corollary 4.3.37].
Lemma 3.4.
Let $H\in\mathbb{C}^{n\times n}$ be Hermitian, and let $\{\mathbf{q}_{k}\}_{k=1}^{m}\subset\mathbb{C}^{n}\,(1\leq m\leq n)$ be a set of orthonormal vectors. Then, for any $i=1,\ldots,m$, it holds that
$\lambda_{i}(H)\leq\lambda_{i}(\check{H})\leq\lambda_{i+n-m}(H),$
where $\check{H}=\big{(}\mathbf{q}_{i}^{\ast}H\mathbf{q}_{j}\big{)}\in\mathbb{C}^{m\times m}$ and $(\cdot)^{\ast}$ denotes the conjugate transpose of a vector.
The following corollary presents a two-sided estimate for $\|E_{\rm TG}\|_{A}$.
Corollary 3.5.
Under the assumptions of Theorem 3.3, it holds that
(3.11) $\sqrt{1-\lambda_{n-r+s+1}(\widetilde{M}A)}\leq\|E_{\rm TG}\|_{A}\leq\sqrt{1-\lambda_{n-r+1}(\widetilde{M}A)}.$
Proof.
Since $I-\Pi$ is an $L^{2}$-orthogonal projector of rank $n-s$, there exists a matrix with orthonormal columns, $V_{n-s}\in\mathbb{R}^{n\times(n-s)}$, such that
$I-\Pi=V_{n-s}V_{n-s}^{T}.$
According to the proof of Theorem 3.3, it follows that $A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}V_{n-s}V_{n-s}^{T}$ has $r-s$ positive eigenvalues and eigenvalue 00 with multiplicity $n-r+s$. Then
$\sigma_{\rm TG}$ $=\lambda_{n-r+s+1}\big{(}\widetilde{M}A(I-\Pi_{A})\big{)}$
$=\lambda_{n-r+s+1}\big{(}A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}(I-\Pi)\big{)}$
$=\lambda_{n-r+s+1}\big{(}A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}V_{n-s}V_{n-s}^{T}\big{)}$
$=\lambda_{n-r+1}\big{(}V_{n-s}^{T}A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}V_{n-s}\big{)}.$
Using Lemma 3.4, we obtain
$\lambda_{n-r+1}(\widetilde{M}A)\leq\sigma_{\rm TG}\leq\lambda_{n-r+s+1}(\widetilde{M}A),$
which, together with (3.7), yields the estimate (3.11). ∎
@@
From the relation
$I-A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}=\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}\big{(}I-A^{\frac{1}{2}}M^{T}A^{\frac{1}{2}}\big{)},$
we deduce that
$\lambda(\widetilde{M}A)=\lambda\big{(}A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}\big{)}\subset(-\infty,1].$
@@
$\lambda(\widetilde{M}A)$ $=\lambda\big{(}I-\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}\big{(}I-A^{\frac{1}{2}}M^{T}A^{\frac{1}{2}}\big{)}\big{)}$
@@
where $\overline{M}$ is defined by (2.5). Hence, $\lambda(\widetilde{M}A)\subset[0,1]$.
@@ 4. An inexact variant of Algorithm 1 and its analysis
In this section, we analyze the convergence of an inexact variant of Algorithm 1, which is described by Algorithm 2. The output of its correction step, $\hat{\mathbf{e}}_{\rm c}$, is expected to be a good approximation to $\mathbf{e}_{\rm c}=A_{\rm c}^{\dagger}\mathbf{r}_{\rm c}$.
@@
3:Coarse-grid correction: $\hat{\mathbf{e}}_{\rm c}\leftarrow\mathscr{B}_{\rm c}\llbracket\mathbf{r}_{\rm c}\rrbracket$triangleright $\mathscr{B}_{\rm c}\llbracket\cdot\rrbracket\approx A_{\rm c}^{\dagger}:\mathbb{R}^{n_{\rm c}}\rightarrow\mathbb{R}^{n_{\rm c}}$

### v2

@@ 3. Convergence analysis of Algorithm 1
Let
$r=\operatorname*{rank}(A)\quad\text{and}\quad s=\operatorname*{rank}(A_{\rm c}).$
@@
which, together with (2.9), yields $\|E_{\rm TG}\|_{A}=0$. However, such a case seldom occurs in practice. In what follows, we only consider the nontrivial case $s<r$.
@@
The convergence factor $\|E_{\rm TG}\|_{A}$ can be reformulated as
@@
where $r=\operatorname*{rank}(A)$ and $F_{\rm TG}$ is given by (2.10).
@@
$=1-\min_{\mathbf{v}_{r}\in\mathbb{R}^{r}\backslash\{0\}}\frac{\mathbf{v}_{r}^{T}\big(I_{r}\,\ 0\big)Q_{1}^{T}F_{\rm TG}Q_{1}\begin{pmatrix}I_{r}\\ 0\end{pmatrix}\mathbf{v}_{r}}{\mathbf{v}_{r}^{T}\big(I_{r}\,\ 0\big)Q_{1}^{T}Q_{1}\begin{pmatrix}I_{r}\\ 0\end{pmatrix}\mathbf{v}_{r}}$
[…]
@@
$=\max_{\mathbf{v}_{r}\in\mathbb{R}^{r}\backslash\{0\}}\frac{\mathbf{v}_{r}^{T}\big(I_{r}\,\ 0\big)Q_{1}^{T}(I-F_{\rm TG})^{2}Q_{1}\begin{pmatrix}I_{r}\\ 0\end{pmatrix}\mathbf{v}_{r}}{\mathbf{v}_{r}^{T}\big(I_{r}\,\ 0\big)Q_{1}^{T}Q_{1}\begin{pmatrix}I_{r}\\ 0\end{pmatrix}\mathbf{v}_{r}}$
[…]
@@
Based on (3.1) and its proof, we can further derive the following identity, which is the main result of this section.
@@
Then, it holds that
(3.6) $\|E_{\rm TG}\|_{A}=\sqrt{1-\lambda_{n-r+s+1}\big(\widetilde{M}A(I-\Pi_{A})\big)},$
where $r=\operatorname*{rank}(A)$ and $s=\operatorname*{rank}(A_{\rm c})$.
@@
$A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}(I-\Pi)+\Pi=Q_{2}\begin{pmatrix}I_{s}&X_{2}\\ 0&X_{3}\end{pmatrix}Q_{2}^{T}.$
Note that $A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}(I-\Pi)+\Pi$ has $r$ positive eigenvalues and nrn-r zero eigenvalues; see the proof of (3.1). We conclude that $X_{3}$ has rsr-s positive eigenvalues and nrn-r zero eigenvalues.
@@
In view of (3.4), we have
$I-A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}=\big(I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big)\big(I-A^{\frac{1}{2}}M^{T}A^{\frac{1}{2}}\big)\succeq 0,$
which leads to
$\lambda_{\max}\big(A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}\big)\leq 1.$
@@
$\lambda\big(A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}\big)$ $=\lambda\big(I-\big(I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big)\big(I-A^{\frac{1}{2}}M^{T}A^{\frac{1}{2}}\big)\big)$
@@
where $\overline{M}$ is defined by (2.5). Hence, $A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}$ is an SPSD matrix with eigenvalues contained in $[0,1]$. Since
$A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}-\big(A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}\big)^{\frac{1}{2}}(I-\Pi)\big(A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}\big)^{\frac{1}{2}}\succeq 0,$
it follows that
$\lambda_{n-r+s+1}\big(\widetilde{M}A(I-\Pi_{A})\big)=\lambda_{n-r+s+1}\big(A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}(I-\Pi)\big)\leq\lambda_{n-r+s+1}(\widetilde{M}A),$
which, combined with (3.6), gives
(3.9) $\|E_{\rm TG}\|_{A}\geq\sqrt{1-\lambda_{n-r+s+1}(\widetilde{M}A)}.$
In the special case where $A\succ 0$ and $\operatorname*{rank}(P)=n_{\rm c}$ (i.e., $r=n$ and $s=n_{\rm c}$), (3.9) reduces to
(3.10) $\|E_{\rm TG}\|_{A}\geq\sqrt{1-\lambda_{n_{\rm c}+1}(\widetilde{M}A)}.$
As proved in [4, Lemma 1], the lower bound in (3.10) is attainable. Furthermore, (3.10) and its sharpness still hold when $A\succeq 0$, $\mathcal{N}(A)\subset\mathcal{R}(P)$, and $\operatorname*{rank}(P)=n_{\rm c}$; see [36, Theorem 5.6].
Using (3.6), we can prove the following corollary.
Corollary 3.5.
Let
$d_{1}=\dim\big(\mathcal{R}(P)\cap\mathcal{N}(A)\big)\quad\text{and}\quad d_{2}=\dim\big(\mathcal{R}(P)+\mathcal{N}(A)\big).$
Then, it holds that
(3.11a) $\|E_{\rm TG}\|_{A}$ $=\sqrt{1-\lambda_{n-r+\operatorname*{rank}(P)-d_{1}+1}\big(\widetilde{M}A(I-\Pi_{A})\big)}$
(3.11b) $=\sqrt{1-\lambda_{d_{2}+1}\big(\widetilde{M}A(I-\Pi_{A})\big)}.$
Proof.
Assume that $\{\mathbf{p}_{1},\ldots,\mathbf{p}_{d_{1}}\}$ is a basis for $\mathcal{R}(P)\cap\mathcal{N}(A)$. Such a set can be expanded as a basis for $\mathcal{R}(P)$, which is denoted by $\big\{\mathbf{p}_{1},\ldots,\mathbf{p}_{d_{1}},\mathbf{p}_{d_{1}+1},\ldots,\mathbf{p}_{\operatorname*{rank}(P)}\big\}$. Let
$P=\big(\mathbf{p}_{1},\ldots,\mathbf{p}_{\operatorname*{rank}(P)}\big)W,$
where $W\in\mathbb{R}^{\operatorname*{rank}(P)\times n_{\rm c}}$ is of full row rank. Then
$AP=\big(0,\ldots,0,A\mathbf{p}_{d_{1}+1},\ldots,A\mathbf{p}_{\operatorname*{rank}(P)}\big)W.$
On the one hand,
$\operatorname*{rank}(AP)\leq\operatorname*{rank}\big(\big(A\mathbf{p}_{d_{1}+1},\ldots,A\mathbf{p}_{\operatorname*{rank}(P)}\big)\big).$
On the other hand,
$\operatorname*{rank}(AP)$ $\geq\operatorname*{rank}\big(\big(A\mathbf{p}_{d_{1}+1},\ldots,A\mathbf{p}_{\operatorname*{rank}(P)}\big)\big)+\operatorname*{rank}(W)-\operatorname*{rank}(P)$
$=\operatorname*{rank}\big(\big(A\mathbf{p}_{d_{1}+1},\ldots,A\mathbf{p}_{\operatorname*{rank}(P)}\big)\big).$
Note that the set $\big\{A\mathbf{p}_{d_{1}+1},\ldots,A\mathbf{p}_{\operatorname*{rank}(P)}\big\}$ is linearly independent. We then have
$\operatorname*{rank}(AP)=\operatorname*{rank}\big(\big(A\mathbf{p}_{d_{1}+1},\ldots,A\mathbf{p}_{\operatorname*{rank}(P)}\big)\big)=\operatorname*{rank}(P)-d_{1}.$
Due to
$\mathcal{N}(A_{\rm c})=\mathcal{N}(A^{\frac{1}{2}}P)=\mathcal{N}(AP),$
it follows that
$s=\operatorname*{rank}(A_{\rm c})=\operatorname*{rank}(AP)=\operatorname*{rank}(P)-d_{1},$
which, together with (3.6), yields the identity (3.11a).
The identity (3.11b) follows from (3.11a) and the fact
$d_{1}=\operatorname*{rank}(P)+n-r-d_{2},$
because
$\dim\big(\mathcal{R}(P)\big)+\dim\big(\mathcal{N}(A)\big)=\dim\big(\mathcal{R}(P)+\mathcal{N}(A)\big)+\dim\big(\mathcal{R}(P)\cap\mathcal{N}(A)\big).$
This completes the proof. ∎
In particular, if $\mathcal{N}(A)\subset\mathcal{R}(P)$, then we have the following result, which provides an alternative to the convergence identity proved in [36, Theorem 5.3].
Corollary 3.6.
If $\mathcal{N}(A)\subset\mathcal{R}(P)$, then
$\|E_{\rm TG}\|_{A}=\sqrt{1-\lambda_{\operatorname*{rank}(P)+1}\big(\widetilde{M}A(I-\Pi_{A})\big)}.$
Proof.
If $\mathcal{N}(A)\subset\mathcal{R}(P)$, then
$d_{1}=n-r\quad\text{and}\quad d_{2}=\operatorname*{rank}(P).$
The desired result then follows immediately from (3.11a) or (3.11b). ∎
@@ 4. An inexact variant of Algorithm 1 and its analysis
In this section, we are devoted to analyzing the convergence of an inexact variant of Algorithm 1 (that is, the coarse-grid system (2.6) is solved approximately), which is described by Algorithm 2 below. The output of its correction step, $\hat{\mathbf{e}}_{\rm c}=\mathscr{B}_{\rm c}\llbracket\mathbf{r}_{\rm c}\rrbracket$, is expected to be a good approximation to $\mathbf{e}_{\rm c}=A_{\rm c}^{\dagger}\mathbf{r}_{\rm c}$. Here, $\mathscr{B}_{\rm c}\llbracket\cdot\rrbracket\approx A_{\rm c}^{\dagger}$ is a general mapping from $\mathbb{R}^{n_{\rm c}}$ to $\mathbb{R}^{n_{\rm c}}$.
@@
3:Coarse-grid correction: $\hat{\mathbf{e}}_{\rm c}\leftarrow\mathscr{B}_{\rm c}\llbracket\mathbf{r}_{\rm c}\rrbracket$